44.0 The paired data below consist of the length of random bears in inches and their weight in pounds. Find and use the correlation equation to predict the weight of a bear that measures 62 inches. 284.0 A 629.765 lbsB 651.765 ibsCy 625
Answer #1

We will find an equation of the regression line in 4 steps. Step 1: Find X Yand X2 as it was done in the table below 584 624 514 700 596 284 550 X. X 3844 4356 1936 4489 3025 1296 2209 62 36208 41184 22616 46900 32780 10224 25850 67 36 47Step 2: Find the sum of every column: ΣΧ= 377, Σs 3852 , ΣΧ.s 215762, ΣΧ2-21155 Step 3: Use the following equations to find a andb Y . X2- X . XY 3852-21 155-377 . 215762 7.21155- 3772 a- 24.645 n . ΣΧΥ-Σχ . ΣΥ 7 . 215762-377 . 3852 9.76 7 21155- (377) Step 4: Substitute a and b in regression equation formula y-24.645 + 9.76.xSo the equation is Y = 24.645 + 9.76 X

Now to find what will be the value of y at X = 62.

Put x =62 in the equation,

We get,

Y = 24.645 + 9.76 × 62

= 629.765 (A) is correct.

Please appreciate it if you like the answer.